OpenStudy (anonymous):

What is the distance between points M and N? A. 14 units B. 8 units C. 48 units D. 10 units E. 6 units

OpenStudy (anonymous):

what cordinates do M and N have?

OpenStudy (anonymous):

exmple: M(3,-2) N(1,0) you should have some information like this ...

OpenStudy (anonymous):

the formula is: \[\LARGE d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2 }\] in MY case you should substitute: \[\LARGE d=\sqrt{(1-3)^2+(0-(-2 ))^2 }\] \[\LARGE d=\sqrt{4+4 }\] \[\LARGE d=2\sqrt{2 }\] so the distance between M(3,-2) N(1,0) would have been \[\LARGE d=2\sqrt{2 }\] hope this helps...

OpenStudy (anonymous):

d--- distance :)